# Unitary Designs of Symmetric Local Random Circuits

### v1

@@ Abstract
We have established the method of characterizing the unitary design generated by a symmetric local random circuit. Concretely, we have shown that the necessary and sufficient condition for the circuit forming an approximate $t$-design is given by simple integer optimization for general symmetry and locality. By using the result, we explicitly give the maximal order of unitary design under the $\mathbb{Z}_{2}$, U(1), and SU(2) symmetries for general locality. This work reveals the relation between the fundamental notions of symmetry and locality in terms of randomness.
Introduction.— Symmetry is one of the most foundational and ubiquitous principles that affect natural phenomena. It not only puts constraints on static or dynamic properties of systems on microscopic and macroscopic scales, but also give rise to new physics as in spontaneous symmetry breaking [1, 2, 3], symmetry-enhanced topological phenomena [4, 5, 6, 7, 8, 9, 10], and quantum memories [11, 12, 13, 14]. The concept has an intriguing interplay with yet another fundamental notion, locality, in both spacial and temporal ways. The former emerges for instance in the hydrodynamic transport [15, 16] and quantum chaos [17, 18], whereas the latter includes the universality of unitary evolution, i.e., the ability to generate arbitrary unitary operations. This is a crucial concept in quantum information science including quantum computing, quantum control, and quantum cryptography.
In the absence of any symmetry, it is well-understood that the universality regarding the global unitary group is achieved by layers of locally universal gate sets [19, 20, 21]. While one may naturally deduce that such a connection between globalness and locality is robust, it was shown that symmetric local gates are not universal for symmetric global gates [22]. This counterintuitive no-go theorem of universality casts light on the impact of symmetry on the practicality of quantum information processing and quantum computing [23], as well as statistical and dynamical properties of quantum many-body systems [24, 25, 26, 27, 28]. An outstanding question we must ask is the following: to what extent the globalness of unitary evolution is allowed under the presence of symmetry for individual quantum operation?
@@
A complementary approach to such a qualitative picture is to examine the expressibility in a quantitative manner via the randomness of local circuits. As known in the non-symmetric cases, this allows us to assess the applicability to quantum tomography [33], quantum device benchmarking [34, 35, 36], variational quantum computation [37], and so on. In this direction, Refs. [38, 39] lower-bounded the expressibility of symmetric circuits with specific locality via the ability to compose symmetric unitary $t$-design, or, mimic the symmetric Haar random unitary distribution up to the $t$th moment  [34, 40, 41]. However, in sharp contrast to the non-symmetric case, it remains an open question to establish a unified theory that provides the maximal order of unitary designs achievable with symmetric local random circuits for general symmetry and locality.
@@
Figure 1: Setup of random circuits and relation between semi-universality and unitary designs with and without symmetry. (a) We consider a symmetric local random circuit in the large depth limit. Local random gates are drawn from a symmetric gate set, which is a set of all local gates commuting with a group representing symmetry, such as $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$. (b) In our setup, semi-universality and unitary designs have rich relation under symmetry, while (c) the relation becomes rather trivial without symmetry.
Setup.— We consider random circuits consisting of symmetric and local gates on $n$ qudits, as shown in Fig 1 (a). We first explain the symmetry constraint and next, introduce the locality constraint. By using the pair of a group $G$ and its unitary representation $R$, we say that an operator is $(G,R)$-symmetric if it commutes with $R(g)$ for all $g\in G$, and we denote the set of all $(G,R)$-symmetric unitary gates on $n$ qudits by $\mathcal{U}_{n,G,R}$. Specifically, the representations of the physically plausible $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries on $n$ qubits are respectively given by
@@
where $\mathrm{X}$, $\mathrm{Y}$, and $\mathrm{Z}$ are the Pauli operators, and the superscript “tot” means the sum of the Pauli operators acting on every single qubit. These symmetries appear in the transverse-field Ising model, the Heisenberg XXZ model, and the XXX model, for example.
Next, by adding the locality constraints, we define symmetric local random circuits. Given $\gamma$ representing a set of labels for qudits, we denote the set of all $(G,R)$-symmetric unitary gates nontrivially acting on the qudit set $\gamma$ by $\mathcal{U}_{n,G,R}^{\gamma}$. We construct a random circuit by taking $\gamma$ from the set $\Gamma$ of qudit sets with nonzero probability $p^{\gamma}>0$ and randomly drawing a unitary gate $U$ from $\mathcal{U}_{n,G,R}^{\gamma}$, i.e., the distribution of unitary gates at each time step is given by
@@
where $\mu_{\mathcal{U}_{n,G,R}^{\gamma}}$ is the normalized Haar measure on $\mathcal{U}_{n,G,R}^{\gamma}$. For example, when we express the nearest neighbor gate sets on a one-dimensional system consisting of $n$ qudits labeled as $1$, $2$, …, $n$ with the open boundary condition, $\Gamma$ is given by $\{\{j,j+1\}|j=1,2,...,n-1\}$, which corresponds to the circuit shown in Fig. 1 (a).
In order to investigate the randomness of symmetric random circuits specified by the distribution $\zeta$, we rely on a notion called approximate unitary design, defined as follows.
@@
(Approximate symmetric unitary design.) Let $\nu$ be a distribution on the unitary group on $n$ qudits. $\nu$ is an approximate $(G,R)$-symmetric unitary $t$-design if
@@
This condition implies that the behavior of the distribution $\nu$ asymptotically coincides with the symmetric global random distribution up to the $t$th order in the infinite depth. We note that even if we assume the infinite depth, the problem of finding the maximal order of $t$ is still highly nontrivial in the presence of symmetry, because there are rich structures in the classes of designs in the presence of symmetry unlike in the absence of symmetry (see Fig. 1 (b)(c)). In the following, we assume that $\Gamma$ is inseparable, i.e., we cannot decompose the qudits into two disjoint parts $C_{1}$ and $C_{2}$ such that every $\gamma\in\Gamma$ satisfies $\gamma\subset C_{1}$ or $\gamma\subset C_{2}$. This is because if $\Gamma$ is separable, the gate set $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$ does not scramble between the two qudit regions, and does not generate unitary designs.
Results for concrete symmetries.— We focus on the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries in our first main result, which explicitly gives the maximal order of approximate unitary design.
@@
($\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries.) Let $n\in\mathbb{N}$, $\Gamma$ be inseparable, $k=\max_{\gamma\in\Gamma}\#\gamma$, $n\geq k+1$, $G=\mathbb{Z}_{2}$, $\mathrm{U}(1)$, or $\mathrm{SU}(2)$, and $R$ be defined by Eq. (1), (2), or (3), respectively. Then, the distributions $\zeta$ defined by Eq. (4) for $(G,R)$-symmetric $k$-local random circuits form approximate $(G,R)$-symmetric unitary $t$-designs if and only if
@@
We note that this result depends only on the locality $k$ of $\Gamma$, and not on other information of $\Gamma$ such as the connectivity between qubits, because every symmetric $k$-local operator can be generated with some symmetric $k$-local operator and swap operators in $\mathcal{U}_{n,G,R}^{\gamma}$ with some $\gamma$ satisfying #$\#\gamma=k$. This result is also independent of the detailed probability $p^{\gamma}$. We present the result for $k=2,3,4$ and the range of $n$ in Table 1. As for the maximal order $t$ for small $n$, there are some exceptions in the $\mathrm{SU}(2)$ case (see Theorem 6 of Ref. [42]). We note that although Eq. (7) holds for all $n\geq k+1$ when $k=2$, $3$ and $4$ in the $\mathrm{U}(1)$ symmetry, this relation does not hold for all $n\geq k+1$.
Table 1: Necessary and sufficient conditions on $t$ for the distribution $\zeta$ of the symmetric local random circuit forming an approximate symmetric unitary $t$-design.
@@
$\mathbb{Z}_{2}$-symmetry $t<2^{n-1}$ ($n\geq 3$) $t<2^{n-1}$ ($n\geq 4$) $t<2^{n-1}$ ($n\geq 5$)
$\mathrm{U}(1)$-symmetry $t<2(n-1)$ ($n\geq 3$) $t<n(n-2)$ ($n\geq 4$) $t<2(n-1)(n-3)$ ($n\geq 5$)
$\mathrm{SU}(2)$-symmetry $t<(n-1)(n-3)$ $(n\geq 9)$ $t<(n-1)(n-3)$ $(n\geq 9)$ $t<2(n-1)(n-3)(n-5)/3$ $(n\geq 13)$
Result for general case.— Next, we present the result for general symmetries. For stating the result, we prepare two notions: semi-universality and symmetric operator decomposition. The semi-universality is a weaker version of universality, defined as follows:
@@
The gate set $\mathcal{X}\subset\mathcal{U}_{n,G,R}$ is semi-universal for $\mathcal{U}_{n,G,R}$ if
@@
We note that this definition means that a semi-universal gate set can generate all operators in $\mathcal{U}_{n,G,R}$ up to $Z(\mathcal{U}_{n,G,R})$, which is the relative phase group between symmetry sectors explained below. When the symmetry is trivial, i.e., $G=\{I\}$, the condition of semi-universality reduces to the ordinary universality up to the global phase, because the center of unitary group $\mathcal{U}_{n}$ is the global phase group $\{e^{i\theta}I\}_{\theta\in\mathbb{R}}$. The semi-universality was first considered in the context of decoherence-free subspaces [29], and it is known that the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, or $\mathrm{SU}(2)$-symmetric $2$-local gate sets satisfy semi-universality but do not satisfy universality [30, 31].
As for the relation to unitary designs, it is known that if the gate set $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ is not semi-universal for $\mathcal{U}_{n,G,R}$, the distribution $\zeta$ defined by Eq. (4) is not an approximate $(G,R)$-symmetric unitary $2$-design, which follows from the combination of Theorem 16 of Ref. [43] (see below Definition 2 in Ref. [42] for details). We note that this statement holds because $\{\mathcal{U}_{n,G,R}^{\gamma}\}_{\gamma\in\Gamma}$ are continuous gate sets, and it does not generally hold for discrete gate sets such as the Clifford gates. On the other hand, the semi-universality does not even imply $\zeta$ being an approximate $1$-design. As a counterexample, for the $\mathbb{Z}_{2}$ symmetry on a single qubit, the trivial gate set $\{I\}$ is semi-universal for $\mathcal{U}_{1,G,R}$, but does not generate a $1$-design.
As the second notion for stating our general result, we introduce the decomposition of symmetric operators. We note that every unitary representation $R$ can be decomposed into inequivalent irreducible representations:
$R(g)\cong\bigoplus_{\lambda\in\Lambda}R_{\lambda}(g)\otimes\mathrm{id}(\mathbb{C}^{m_{\lambda}}),$ (10)
where $\Lambda$ is the set of all labels $\lambda$ for the irreducible representations $R_{\lambda}$ appearing in the decomposition of $R$, $m_{\lambda}$ is the multiplicity of $R_{\lambda}$, $\mathrm{id}(\mathbb{C}^{m})$ is the identity operator on $\mathbb{C}^{m}$, and $\cong$ means isomorphism. Then, by Schur’s lemma, every $(G,R)$-symmetric operator $A$ can be shown to be written as [44]
@@
with some uniquely determined operator $A_{\lambda}$ acting on the multiplicity space $\mathbb{C}^{m_{\lambda}}$, where $r_{\lambda}$ is the dimension of the representation space of $R_{\lambda}$. By using this decomposition, we can confirm that $Z(\mathcal{U}_{n,G,R})\cong\bigoplus_{\lambda\in\Lambda}e^{i\theta_{\lambda}}\mathrm{id}(\mathbb{C}^{r_{\lambda}})\otimes\mathrm{id}(\mathbb{C}^{m_{\lambda}})$ is the relative phase group, and that the semi-universality of $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ for $\mathcal{U}_{n,G,R}$ is equivalent to the following condition: every unitary operator in the form of $\bigoplus_{\lambda\in\Lambda}\mathrm{id}(\mathbb{C}^{r_{\lambda}})\otimes U_{\lambda}$ with $U_{\lambda}\in\mathrm{SU}(m_{\lambda})$ is generated by $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$.
We illustrate the decomposition by taking the $\mathrm{U}(1)$ symmetry. Since $\mathrm{U}(1)$ is commutative, every irreducible representation is one-dimensional and corresponds to the eigenvalue of $\mathrm{Z}^{\mathrm{tot}}$. Concretely, it is decomposed into $n+1$ irreducible representations $R_{\lambda}(e^{i\theta})=e^{i(n-2\lambda)\theta}$ with $\lambda\in\{0,1,...,n\}$, and the multiplicity $m_{\lambda}$ is given by $\binom{n}{\lambda}$, which corresponds to the dimension of the eigenspace of $\mathrm{Z}^{\mathrm{tot}}$ with eigenvalue $n-2\lambda$.
@@
(General case.) Let $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ be semi-universal for $\mathcal{U}_{n,G,R}$. Then, $\zeta$ is an approximate $(G,R)$-symmetric unitary $t$-design if and only if
$t<\min_{\bm{x}\in(\mathcal{V}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}}(\bm{m},\bm{x}^{+}),$ (12)
@@
where $\mathcal{V}:=\mathrm{span}(\{\bm{f}(A)|A\in\mathcal{L}_{n,G,R}^{\gamma},\gamma\in\Gamma\})$, $\mathcal{L}_{n,G,R}^{\gamma}$ is the set of all $(G,R)$-symmetric linear operators nontrivially acting on the qudits labeled by $\gamma$, $\bm{f}(A)=(f_{\lambda}(A))_{\lambda\in\Lambda}:=(\mathrm{tr}(A_{\lambda}))_{\lambda\in\Lambda}$ with $A_{\lambda}$ in Eq. (11) for symmetric operators $A$, $\bm{m}:=(m_{\lambda})_{\lambda\in\Lambda}$, $\bm{x}^{+}:=((|x_{\lambda}|+x_{\lambda})/2)_{\lambda\in\Lambda}$, and we use the standard inner product in $\mathbb{C}^{\Lambda}$.

### v3

@@ Abstract
We have established the method of characterizing the unitary design generated by a symmetric local random circuit. Concretely, we have shown that the necessary and sufficient condition for the circuit asymptotically forming a $t$-design is given by simple integer optimization for general symmetry and locality. By using the result, we explicitly give the maximal order of unitary design under the $\mathbb{Z}_{2}$, U(1), and SU(2) symmetries for general locality. This work reveals the relation between the fundamental notions of symmetry and locality in terms of randomness.
Introduction.— Symmetry is one of the most foundational and ubiquitous principles that affect natural phenomena. It not only puts constraints on static or dynamic properties of systems on microscopic and macroscopic scales, but also gives rise to new phenomena and techniques such as spontaneous symmetry breaking [1, 2, 3], symmetry-enhanced topological phenomena [4, 5, 6, 7, 8, 9, 10], quantum memories [11, 12, 13, 14], decoherence-free subspaces [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], covariant quantum error-correcting codes [26, 27, 28, 29, 30, 31, 32, 33, 34], and geometric quantum machine learning [35, 36, 37, 38, 39, 40, 41, 42, 43, 44]. The concept has an intriguing interplay with yet another fundamental notion, locality, in both spatial and temporal ways. The former emerges for instance in the hydrodynamic transport [45, 46] and quantum chaos [47, 48], whereas the latter includes the universality of unitary evolution, i.e., the ability to generate arbitrary unitary operations. This is a crucial concept in quantum information science including quantum computing, quantum control, and quantum cryptography.
In the absence of any symmetry, it is well understood that the universality regarding the global unitary group is achieved by layers of locally universal gate sets [49, 50, 51]. While one may naturally deduce that such a connection between globalness and locality is robust, it was shown that symmetric local gates are not universal for symmetric global gates [52]. This counterintuitive no-go theorem of universality casts light on the impact of symmetry on the practicality of quantum information processing and quantum computing [53], as well as the statistical and dynamical properties of quantum many-body systems [54, 55, 56, 57, 58, 59, 60]. An outstanding question we must ask is the following: to what extent the globalness of unitary evolution is allowed under the presence of symmetry for individual quantum operations?
@@
A complementary approach to such a qualitative picture is to examine the expressibility in a quantitative manner via the randomness of local circuits. As known in the non-symmetric cases, this allows us to assess the applicability to quantum tomography [64], quantum device benchmarking [65, 66, 67], variational quantum computation [68], and so on. In this direction, Refs. [69, 70] lower-bounded the expressibility of symmetric circuits with specific locality via the ability to compose a symmetric unitary $t$-design, or, mimic the symmetric Haar random unitary distribution up to the $t$th moment [65, 71, 72]. However, in sharp contrast to the non-symmetric case, it remains an open question to establish a unified theory that provides the maximal order of unitary designs achievable with symmetric local random circuits for general symmetry and locality.
@@
Figure 1: Setup of random circuits and relation between semi-universality and unitary designs with and without symmetry. (a) We consider a symmetric local random circuit in the large depth limit. Local random gates are drawn from a symmetric gate set, which is a set of all local gates commuting with a group representing symmetry, such as $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$. (b) In our setup, semi-universality (see Definition 2 for explanation) and unitary designs have rich relation under symmetry, while (c) the relation becomes rather trivial without symmetry.
Setup.— We consider random circuits consisting of symmetric and local gates on $n$ qudits, as shown in Fig. 1 (a). We first explain the symmetry constraint and next, introduce the locality constraint. By using the pair of a group $G$ and its unitary representation $R$, we say that an operator is $(G,R)$-symmetric if it commutes with $R(g)$ for all $g\in G$, and we denote the set of all $(G,R)$-symmetric unitary gates on $n$ qudits by $\mathcal{U}_{n,G,R}$. Specifically, the representations of the physically plausible $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries on $n$ qubits are respectively given by
@@
where $\mathrm{X}$, $\mathrm{Y}$, and $\mathrm{Z}$ are the Pauli operators, and the superscript “tot” means the sum of the Pauli operators acting on every single qubit. These symmetries appear in the transverse-field Ising model, the Heisenberg XXZ model, and the XXX model, respectively, for example.
Next, by adding the locality constraints, we define symmetric local random circuits. Given $\gamma$ representing a set of labels for qudits, we denote the set of all $(G,R)$-symmetric unitary gates nontrivially acting on the qudit set $\gamma$ by $\mathcal{U}_{n,G,R}^{\gamma}$. We construct a random circuit by taking $\gamma$ from the set $\Gamma$ of qudit sets with nonzero probability $p^{\gamma}>0$ and randomly drawing a unitary gate $U$ from $\mathcal{U}_{n,G,R}^{\gamma}$. Then, the distribution of unitary gates at each time step is given by
@@
where $\mu_{\mathcal{U}_{n,G,R}^{\gamma}}$ is the normalized Haar measure on $\mathcal{U}_{n,G,R}^{\gamma}$. For example, when we express the nearest neighbor gate sets on a $1$-dimensional system consisting of $n$ qudits labeled as $1$, $2$, …, $n$ with the open boundary condition, $\Gamma$ is given by $\{\{j,j+1\}|j=1,2,...,n-1\}$, which corresponds to the circuit shown in Fig. 1 (a). We define the locality $k$ of $\Gamma$ as the maximal number of qudits that we can simultaneously manipulate, i.e., $k:=\max_{\gamma\in\Gamma}\#\gamma$, where #$\#\gamma$ means the size of a set $\gamma$. In the following, in order to exclude trivial cases, we assume that the circuit cannot be decomposed into any two independent parts, i.e., if we decompose the qudits into any two parts, there exists some gate set $\mathcal{U}_{n,G,R}^{\gamma}$ that acts on both parts.
In order to characterize the randomness of the symmetric random circuits specified by the distribution $\zeta$, we rely on a notion called asymptotic unitary design, defined as follows.
@@
(Asymptotic symmetric unitary design.) Let $\nu$ be a distribution on the unitary group on $n$ qudits. $\nu$ is an asymptotic $(G,R)$-symmetric unitary $t$-design if
@@
This condition implies that the behavior of the distribution $\nu$ asymptotically coincides with the symmetric global Haar random distribution up to the $t$th order in the infinite depth. Our definition of asymptotic unitary design is equivalent to forming an $\epsilon$-approximate unitary design for all $\epsilon>0$ by taking a sufficiently large depth $D$. We note that even if we assume the infinite depth, the problem of finding the maximal order of $t$ is still highly nontrivial in the presence of symmetry as opposed to the non-symmetric case, where there does not exist an upper bound on the achievable order $t$ [73].
Results for concrete symmetries.— We focus on the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, and $\mathrm{SU}(2)$ symmetries in our first main result, which explicitly gives the tight bound on the maximal achievable order of asymptotic unitary design for sufficiently large $n$.
@@
(Concrete symmetries.) Let $n\geq k+1$, $G=\mathbb{Z}_{2}$, $\mathrm{U}(1)$, or $\mathrm{SU}(2)$, and $R$ be defined by Eq. (1), (2), or (3), respectively. Then, the distributions $\zeta$ defined by Eq. (4) for $(G,R)$-symmetric $k$-local random circuits form asymptotic $(G,R)$-symmetric unitary $t$-designs if and only if
@@
We note that this result depends only on the locality $k$ of $\Gamma$, and not on other information about $\Gamma$ such as the connectivity between qubits, or the detailed probability $p^{\gamma}$, because every symmetric $k$-local unitary can be written as the product of some unitary $U\in\mathcal{U}_{n,G,R}^{\gamma}$ satisfying #$\#\gamma=k$ and some permutation operator, which are also generated by products of swap operators included in some $\mathcal{U}_{n,G,R}^{\gamma}$. In Table 1, for the cases of $k=2$, $3$, and $4$, we present the ranges of the qubit count $n$ for which Eqs. (7) and (8) give the tight upper bounds on the achievable orders of unitary designs (see Theorems 4 and 6 of Ref. [74] for smaller qubit count $n$). Although Eq. (7) gives the tight bound for all $n\geq k+1$ when $k=2$, $3$ and $4$, it is not necessarily tight when $k\geq 5$. We finally note that even when $n$ is not sufficiently large and Eqs. (7) and (8) do not give the tight bound, they give upper bounds on the achievable orders of unitary designs for any $n\geq k+1$.
Table 1: Necessary and sufficient conditions on $t$ for the distribution $\zeta$ of the symmetric local random circuit forming an asymptotic symmetric unitary $t$-design.
@@
$\mathbb{Z}_{2}$ symmetry $t<2^{n-1}$ ($n\geq 3$) $t<2^{n-1}$ ($n\geq 4$) $t<2^{n-1}$ ($n\geq 5$)
$\mathrm{U}(1)$ symmetry $t<2(n-1)$ ($n\geq 3$) $t<n(n-2)$ ($n\geq 4$) $t<2(n-1)(n-3)$ ($n\geq 5$)
$\mathrm{SU}(2)$ symmetry $t<(n-1)(n-3)$ $(n\geq 9)$ $t<(n-1)(n-3)$ $(n\geq 9)$ $t<2(n-1)(n-3)(n-5)/3$ $(n\geq 13)$
Result for general case.— We present the result for general symmetries. For stating the result, we prepare two notions: semi-universality and symmetric operator decomposition. While a unitary gate set is called universal for $\mathcal{U}_{n,G,R}$ if it can generate $\mathcal{U}_{n,G,R}$, the semi-universality is a more relaxed condition defined as follows:
@@
(Semi-universality.) The gate set $\mathcal{X}\subset\mathcal{U}_{n,G,R}$ is semi-universal for $\mathcal{U}_{n,G,R}$ if
@@
This definition means that a semi-universal gate set can generate all operators in $\mathcal{U}_{n,G,R}$ up to the freedom of $Z(\mathcal{U}_{n,G,R})$, which is the relative phase group between symmetry sectors explicitly shown below as Eq. (12). In the absence of symmetry, i.e., $G=\{I\}$, the condition of semi-universality reduces to the ordinary universality up to the global phase, because the center of unitary group $\mathcal{U}_{n}$ is the global phase group $\{e^{i\theta}I\}_{\theta\in\mathbb{R}}$. The semi-universality was introduced in the context of decoherence-free subspaces [25], and it is known that the $\mathbb{Z}_{2}$, $\mathrm{U}(1)$, or $\mathrm{SU}(2)$-symmetric $2$-local gate sets are not universal but semi-universal [61, 62].
We clarify the relation between (semi-)universality and unitary design shown in Figs. 1 (b) and (c). First, if the distribution $\zeta$ defined by Eq. (4) is an asymptotic $(G,R)$-symmetric unitary $2$-design, the gate set $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ is semi-universal for $\mathcal{U}_{n,G,R}$, which follows from Theorem 16 of Ref. [75] (see below Definition 2 in Ref. [74] for details). The converse is not true; As a counterexample, under the $\mathbb{Z}_{2}$ symmetry on a single qubit, the trivial gate set $\{I\}$ is semi-universal for $\mathcal{U}_{1,G,R}$, but does not generate even a $1$-design. Second, the universality of $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ for $\mathcal{U}_{n,G,R}$ is equivalent to $\zeta$ forming an asymptotic unitary $t$-design for all $t\in\mathbb{N}$, as we explain below Theorem 2. This implies that in the non-symmetric case, the semi-universality and the universality coincide, and thus all the classes of unitary $t$-designs with $t\geq 2$ collapse into $\infty$-designs. We note that this statement holds only for continuous gate sets, and not for discrete gate sets such as the Clifford gates.
Next, we introduce the decomposition of symmetric operators. Let us start from the fact that every unitary representation $R$ can be decomposed into inequivalent irreducible representations:
$R(g)\cong\bigoplus_{\lambda\in\Lambda}R_{\lambda}(g)\otimes\mathrm{id}(\mathbb{C}^{m_{\lambda}})\ \forall g\in G,$ (10)
where $\Lambda$ is the set of all labels $\lambda$ for the irreducible representations $R_{\lambda}$’s that appear in the decomposition of $R$ with multiplicity $m_{\lambda}$, $\mathrm{id}(\mathbb{C}^{m})$ is the identity operator on $\mathbb{C}^{m}$, and $\cong$ means isomorphism. Then, by Schur’s lemma, every $(G,R)$-symmetric operator $A$ can be written as [76]
@@
with some operators $A_{\lambda}$ acting on the multiplicity spaces $\mathbb{C}^{m_{\lambda}}$, where $r_{\lambda}$ is the dimension of the representation space of $R_{\lambda}$, and $A_{\lambda}$ is uniquely determined for $A$. By using this decomposition, we can confirm that $Z(\mathcal{U}_{n,G,R})$ is the relative phase group:
$Z(\mathcal{U}_{n,G,R})\cong\left\{\bigoplus_{\lambda\in\Lambda}e^{i\theta_{\lambda}}\mathrm{id}(\mathbb{C}^{r_{\lambda}}\otimes\mathbb{C}^{m_{\lambda}})\ \middle|\ \theta_{\lambda}\in\mathbb{R}\right\}.$ (12)
We illustrate the decomposition by taking the $\mathrm{U}(1)$ symmetry. Since $\mathrm{U}(1)$ is commutative, all irreducible representations are $1$-dimensional and correspond to the eigenvalues of $\mathrm{Z}^{\mathrm{tot}}$. Concretely, it is decomposed into $n+1$ irreducible representations $R_{\lambda}(e^{i\theta})=e^{i(n-2\lambda)\theta}$ with $\lambda\in\{0,1,...,n\}$, and its multiplicity $m_{\lambda}$ is given by $\binom{n}{\lambda}$, which corresponds to the dimension of the eigenspace of $\mathrm{Z}^{\mathrm{tot}}$ with eigenvalue $n-2\lambda$.
@@
(General symmetry.) Let $\bigcup_{\gamma\in\Gamma}\mathcal{U}_{n,G,R}^{\gamma}$ be semi-universal for $\mathcal{U}_{n,G,R}$. Then, the distribution $\zeta$ defined by Eq. (4) is an asymptotic $(G,R)$-symmetric unitary $t$-design if and only if
$t<\min_{\bm{x}\in(\mathcal{C}^{\perp}\cap\mathbb{Z}^{\Lambda})\backslash\{\bm{0}\}}\braket{\bm{m},\bm{x}^{+}},$ (13)
@@
where $\mathcal{C}:=\mathrm{span}(\{\bm{f}(A)|A\in\mathcal{L}_{n,G,R}^{\gamma},\gamma\in\Gamma\})$, $\mathcal{L}_{n,G,R}^{\gamma}$ is the set of all $(G,R)$-symmetric linear operators nontrivially acting on the qudits labeled by $\gamma$, $\bm{f}(A)=(f_{\lambda}(A))_{\lambda\in\Lambda}:=(\mathrm{tr}(A_{\lambda}))_{\lambda\in\Lambda}$ with $A_{\lambda}$ in Eq. (11) for symmetric operators $A$, $\bm{m}:=(m_{\lambda})_{\lambda\in\Lambda}$, $\bm{x}^{+}=(x_{\lambda}^{+})_{\lambda\in\Lambda}:=((|x_{\lambda}|+x_{\lambda})/2)_{\lambda\in\Lambda}$, we use the standard inner product $\braket{\bm{a},\bm{b}}:=\sum_{\lambda\in\Lambda}a_{\lambda}^{*}b_{\lambda}$ for $\bm{a},\bm{b}\in\mathbb{C}^{\Lambda}$, and $\mathcal{C}^{\perp}$ is the orthogonal complement of $\mathcal{C}$ with respect to the inner product.